# Point-hyperplane incidences via extremal graph theory

### v1

@@ 5.2 Constructions with no large complete bipartite subgraphs
In this section, we prove the upper bounds of Theorem 1.3. However, instead of assuming $m\geq n$ and giving bounds on both $\operatorname{rs}_{d}(m,n,I)$ and $\operatorname{rs}_{d}(n,m,I)$, as originally stated in Theorem 1.3, we prefer to keep the notation consistent by always denoting the number of hyperplanes by $m$ and the number of points by $n$. Thus, we will give bounds on $\operatorname{rs}_{d}(m,n,I)$ which depend on $\max\{m,n\}$ and $\min\{m,n\}$. With this notational change, Theorem 1.3 can be equivalently restated as follows.
@@
Let $\varepsilon\in(0,1)$, and let $d$ and $n,m$ be integers which satisfy
@@
Then there exists a prime $p$, a set $\mathcal{P}$ of $n$ points and a set $\mathcal{H}$ of $m$ hyperplanes in $\mathbb{F}_{p}^{d}$ such that $I(\mathcal{P},\mathcal{H})\geq\varepsilon mn$ and $\operatorname{rs}(\mathcal{P},\mathcal{H})\leq(2s)^{4}\varepsilon\max\{m,n\}$, where $s$ is given by Lemma 5.6.
@@
For notational convenience, we assume that $\max\{m,n\}=m$, noting that the same argument applies if $\max\{m,n\}=n$. Let $p$ be a prime satisfying $\varepsilon^{-1}/2<p\leq\varepsilon^{-1}$, and set $t=\lfloor\log_{p}4m\rfloor$ and $\ell=\lceil\frac{4m}{p^{t}}\rceil\leq p$. Note that $4m\leq\ell p^{t}\leq 8m$. Since $p>\varepsilon^{-1}/2\geq 2m^{1/d}$, we have $p^{d}>4m$ and so $t\leq d-1$. Thus, Lemma 5.6 can be applied with $k=\ell$ to get a set of points $\mathcal{P}_{0}$ and a set of hyperplanes $\mathcal{H}_{0}$ with $n_{0}=|\mathcal{P}_{0}|\geq\ell p^{t}/2$, $m_{0}=|\mathcal{H}_{0}|\geq\ell p^{t}/4$ and $I(\mathcal{P}_{0},\mathcal{H}_{0})=\frac{1}{p}m_{0}n_{0}$, satisfying (i), (ii) and (iii).
Let us now argue that $\operatorname{rs}(\mathcal{P}_{0},\mathcal{H}_{0})\leq(2s)^{4}\varepsilon m$. Let $A\subseteq\mathcal{P}_{0},B\subseteq\mathcal{H}_{0}$ such that every point of $A$ lies on every hyperplane in $B$. We consider three cases: either $|A|\geq s^{2}\ell p^{t-2}$, $|B|\geq s^{2}\ell p^{t-2}$ or $\max\{|A|,|B|\}<s^{2}\ell p^{t-2}$. In all cases, we show that
$|A|\cdot|B|\leq s^{4}\ell p^{t-1}\leq 2s^{4}\varepsilon\ell p^{t}\leq 16s^{4}\varepsilon m.$
@@
To finish the proof, let $\mathcal{P}$ be a random subset of $\mathcal{P}_{0}$ containing $n$ points and $\mathcal{H}$ be a random subset of $\mathcal{H}_{0}$ containing $m$ hyperplanes. One should verify that $n\leq|\mathcal{P}_{0}|$ and $m\leq|\mathcal{H}_{0}|$, but this is clear since $|\mathcal{P}_{0}|,|\mathcal{H}_{0}|\geq\frac{1}{4}\ell p^{t}\geq m$ and $m\geq n$. We also have $\operatorname{rs}(\mathcal{P},\mathcal{H})\leq\operatorname{rs}(\mathcal{P}_{0},\mathcal{H}_{0})\leq(2s)^{4}\varepsilon m$. Furthermore, the expected number of incidences between $\mathcal{P}$ and $\mathcal{H}$ is
$\mathbb{E}\big{[}I(P,\mathcal{H})\big{]}=\frac{n}{n_{0}}\frac{m}{m_{0}}I(P_{0},\mathcal{H}_{0})=\frac{1}{p}mn\geq\varepsilon mn.$
Thus, there exist subsets $\mathcal{P},\mathcal{H}$ of $\mathcal{P}_{0},\mathcal{H}_{0}$ with at least $\varepsilon mn$ incidences and $\operatorname{rs}(P,\mathcal{H})\leq(2s)^{4}\varepsilon m$. ∎
@@
Let $\varepsilon\in(0,1)$, and let $d$ and $n\leq m$ be positive integers which satisfy
$\varepsilon\leq\frac{1}{4}n^{-\frac{1}{d-1}}.$
Then there exists a prime power $q$, a set $\mathcal{P}$ of $n$ points and a set $\mathcal{H}$ of $m$ hyperplanes in $\mathbb{F}_{q}^{d}$ such that $I(\mathcal{P},\mathcal{H})\geq\varepsilon mn$ and $\operatorname{rs}(\mathcal{P},\mathcal{H})\leq(4s)^{4}\varepsilon m$, where $s$ is given by Lemma 5.6.
@@
More precisely, we apply Lemma 5.7 to find a set of $n$ points $\mathcal{P}_{0}$ and a set of $n$ hyperplanes $\mathcal{H}_{0}$ in $\mathbb{F}_{p}^{d-1}$ for some prime $p$ for which $I(\mathcal{P}_{0},\mathcal{H}_{0})\geq\varepsilon n^{2}$ and $\operatorname{rs}(\mathcal{P}_{0},\mathcal{H}_{0})\leq(2s)^{4}\varepsilon n$. The hyperplanes and points may be considered in the larger ambient space $\mathbb{F}_{q}^{d-1}$, for any prime power $q=p^{k}$. Passing to $\mathbb{F}_{q}^{d-1}$ does not change the incidence graph. In the rest of the proof, we fix $q$ to be a power of $p$ larger than $m/n$.
Now, for each $x\in\mathbb{F}_{q}^{d-1}$, let $x^{\prime}\in\mathbb{F}_{q}^{d}$ be the point whose first $d-1$ coordinates are the same as $x$, and the last coordinate is 00. Set $\mathcal{P}=\{x^{\prime}|x\in\mathcal{P}_{0}\}$. Furthermore, we fix a set $S=\{s_{1},\dots,s_{\ell}\}\subseteq\mathbb{F}_{q}$ of size $\ell=\lceil m/n\rceil$. Then, for every hyperplane $H\in\mathcal{H}_{0}$, if $H$ is given by the equation $\langle a,x\rangle=a_{1}x_{1}+\cdots+a_{d-1}x_{d-1}=b$, we define hyperplanes $H^{(1)},\dots,H^{(\ell)}$ in $\mathbb{F}_{q}^{d}$, where $H^{(i)}$ is given by the equations $a_{1}x_{1}+\cdots+a_{d-1}x_{d-1}+s_{i}x_{d}=b$. Then, we let $\mathcal{H}_{1}=\{H^{(i)}|i\in[\ell],H\in\mathcal{H}_{0}\}$.
The sizes of the newly constructed sets are $|\mathcal{P}|=n,|\mathcal{H}_{1}|=n\lceil\frac{m}{n}\rceil>m$. Furthermore, if $x\in\mathcal{P}_{0}$ is incident to $H\in\mathcal{H}_{0}$, then $x^{\prime}$ is incident to all of the hyperplanes $H^{(1)},\dots,H^{(\ell)}$. Thus, the number of incidences between $\mathcal{P}$ and $\mathcal{H}_{1}$ is $I(\mathcal{P},\mathcal{H}_{1})\geq\ell I(\mathcal{P}_{0},\mathcal{H}_{0})\geq\varepsilon\ell|\mathcal{P}_{0}||\mathcal{H}_{0}|=\varepsilon|\mathcal{P}||\mathcal{H}_{1}|$. Finally, the size of the largest complete bipartite graph is bounded by
$\operatorname{rs}(\mathcal{P},\mathcal{H}_{1})\leq\ell\cdot\operatorname{rs}(\mathcal{P}_{0},\mathcal{H}_{0})\leq\ell\cdot(2s)^{4}\varepsilon n\leq(4s)^{4}\varepsilon m.$
@@
Let $\varepsilon\in(0,1)$, and let $d$ and $n\geq m$ be positive integers which satisfy
$\varepsilon\leq\frac{1}{4}m^{-\frac{1}{d-1}}.$
Then there exists a prime power $q$, a set $\mathcal{P}$ of $n$ points and a set $\mathcal{H}$ of $m$ hyperplanes in $\mathbb{F}_{q}^{d}$ such that $I(\mathcal{P},\mathcal{H})\geq\varepsilon mn$ and $\operatorname{rs}(\mathcal{P},\mathcal{H})\leq(4s)^{4}\varepsilon n$, where $s$ is given by Lemma 5.6.
@@
The proof is verbatim the same as the proof of Lemma 5.8, the only difference being that we “blow up” points instead of hyperplanes. More precisely, we start from a set of $m$ points $\mathcal{P}_{0}$ and a set of $m$ hyperplanes $\mathcal{H}_{0}$ in $\mathbb{F}_{p}^{d-1}$ for which $I(\mathcal{P}_{0},\mathcal{H}_{0})\geq\varepsilon m^{2}$ and $\operatorname{rs}(\mathcal{P}_{0},\mathcal{H}_{0})\leq(2s)^{4}\varepsilon m$ and consider them in $\mathbb{F}_{q}^{d-1}$, where $q>n/m$.
Then, we fix a set $S=\{s_{1},\dots,s_{\ell}\}\subseteq\mathbb{F}_{q}$ of size $\ell=\lceil n/m\rceil$ and for each point $x\in\mathbb{F}_{q}^{d-1}$ we define $x^{(i)}\in\mathbb{F}_{q}^{d}$ be the point whose first $d-1$ coordinates are the same as $x$, and the last coordinate is $s_{i}$. Set $\mathcal{P}_{1}=\{x^{(i)}|x\in\mathcal{P}_{0},i\in[\ell]\}$. Then, for every hyperplane $H\in\mathcal{H}_{0}$, if $H$ is given by the equation $\langle a,x\rangle=a_{1}x_{1}+\cdots+a_{d-1}x_{d-1}=b$, we define the hyperplane $H^{\prime}$ in $\mathbb{F}_{q}^{d}$, given by the equation $a_{1}x_{1}+\cdots+a_{d-1}x_{d-1}+0\cdot x_{d}=b$. Then, we let =$\mathcal{H}=\{H^{\prime}|H\in\mathcal{H}_{0}\}$.
The sizes of the newly constructed sets are $|\mathcal{H}|=m,|\mathcal{P}_{1}|=m\lceil\frac{n}{m}\rceil>n$. Thus, by choosing a random subset $\mathcal{P}\subseteq\mathcal{P}_{1}$ and performing the same calculation as in the proof of Lemma 5.8 one arrives at the desired conclusion. ∎
@@
Let $\varepsilon\in(0,1)$, and let $d$ and $n,m$ be positive integers which satisfy
$4\varepsilon\geq m^{-\frac{1}{d-2}}\text{ and }4\varepsilon\geq n^{-\frac{1}{d-2}}.$
Then there exists a prime power $q$, a set $\mathcal{P}$ of $n$ points and a set $\mathcal{H}$ of $m$ hyperplanes in $\mathbb{F}_{q}^{d}$ such that $I(\mathcal{P},\mathcal{H})\geq\varepsilon mn$ and $\operatorname{rs}(\mathcal{P},\mathcal{H})\leq 2^{6}s^{4}(4\varepsilon)^{d-1}mn$, where $s$ is given by Lemma 5.6.
@@
The proof of this lemma is somewhat similar to the proofs of Lemmas 5.8 and 5.9, since we use the idea of “blowing-up” a lower-dimensional construction. However, in this case we use a $(d-2)$-dimensional configuration as a starting point and we “blow-up” both points and hyperplanes.
@@
As in the proof of Lemma 5.8, we consider $\mathcal{P}_{0},\mathcal{H}_{0}$ in $\mathbb{F}_{q}^{d-2}$ instead of $\mathbb{F}_{p}^{d-2}$, where $q$ is the power of $p$ satisfying $q>\max\{\frac{m}{N},\frac{n}{N}\}$. Furthermore, we fix two sets $S,T\subseteq\mathbb{F}_{q}$, such that $S=\{s_{1},\dots,s_{k}\}$ and $T=\{t_{1},\dots,t_{\ell}\}$, where $k=\lceil\frac{n}{N}\rceil,\ell=\lceil\frac{m}{N}\rceil$. To every point $x\in\mathcal{P}_{0}$, we associate $k$ points of $\mathbb{F}_{q}^{d}$ as follows. For $i\in[k]$, let $x^{(i)}\in\mathbb{F}_{q}^{d}$ be equal to $x$ on the first $d-2$ coordinates, equal to $s_{i}$ on the coordinate $d-1$, and equal to 00 on the last coordinate. Similarly, for a hyperplane $H\in\mathcal{H}_{0}$ given by the equation $a_{1}x_{1}+\cdots+a_{d-2}x_{d-2}=b$, we associate the hyperplanes $H^{(j)}$ in $\mathbb{F}_{q}^{d}$ for $j\in[\ell]$ defined by the equations $a_{1}x_{1}+\cdots+a_{d-2}x_{d-2}+0\cdot x_{d-1}+t_{j}x_{d}=b$. Finally, we set $\mathcal{P}_{1}=\{x^{(i)}|i\in[k],x\in\mathcal{P}_{0}\}$ and $\mathcal{H}_{1}=\{H^{(j)}|H\in\mathcal{H}_{0},j\in[\ell]\}$.
@@
Recalling that $N=(4\varepsilon)^{-(d-2)}$, we obtain $\operatorname{rs}(\mathcal{P}_{1},\mathcal{H}_{1})\leq 2^{6}s^{4}(4\varepsilon)^{d-1}mn$. Subsampling a random $n$ element subset $\mathcal{P}$ of $\mathcal{P}_{1}$, and a random $m$ element subset $\mathcal{H}$ of $\mathcal{H}_{1}$, completes the proof. ∎

### v2

@@ 3.4 Constructions of dense incidence graphs with no large bipartite subgraphs
In this section, we prove the upper bounds of Theorem 1.3. However, instead of assuming $m\leq n$ and giving bounds on both $\operatorname{rs}_{d}(m,n,I)$ and $\operatorname{rs}_{d}(n,m,I)$, as originally stated in Theorem 1.3, we prefer to keep the notation consistent by always denoting the number of hyperplanes by $m$ and the number of points by $n$. Thus, we will give bounds on $\operatorname{rs}_{d}(m,n,I)$ which depend on $\max\{m,n\}$ and $\min\{m,n\}$. With this notational change, Theorem 1.3 can be equivalently restated as follows.
@@
Let $\varepsilon\in(0,1)$, and let $d$ and $m,n$ be integers which satisfy
@@
Then there exists a prime $p$, a set $\mathcal{P}$ of $m$ points and a set $\mathcal{H}$ of $n$ hyperplanes in $\mathbb{F}_{p}^{d}$ such that $I(\mathcal{P},\mathcal{H})\geq\varepsilon mn$ and $\operatorname{rs}(\mathcal{P},\mathcal{H})\leq(2s)^{4}\varepsilon\max\{m,n\}$, where $s$ is given by Lemma 3.14.
@@
For notational convenience, we assume that $\max\{m,n\}=n$, noting that the same argument applies if $\max\{m,n\}=m$. Let $p$ be a prime satisfying $\varepsilon^{-1}/2<p\leq\varepsilon^{-1}$, and set $t=\lfloor\log_{p}4n\rfloor$ and $\ell=\lceil\frac{4n}{p^{t}}\rceil\leq p$. Note that $4n\leq\ell p^{t}\leq 8n$. Since $p>\varepsilon^{-1}/2\geq 2n^{1/d}$, we have $p^{d}>4n$ and so $t\leq d-1$. Thus, Lemma 3.14 can be applied with $k=\ell$ to get a set of points $\mathcal{P}_{0}$ and a set of hyperplanes $\mathcal{H}_{0}$ with $m_{0}=|\mathcal{P}_{0}|\geq\ell p^{t}/2$, $n_{0}=|\mathcal{H}_{0}|\geq\ell p^{t}/4$ and $I(\mathcal{P}_{0},\mathcal{H}_{0})=\frac{1}{p}m_{0}n_{0}$, satisfying (i), (ii) and (iii).
Let us now argue that $\operatorname{rs}(\mathcal{P}_{0},\mathcal{H}_{0})\leq(2s)^{4}\varepsilon n$. Let $A\subseteq\mathcal{P}_{0},B\subseteq\mathcal{H}_{0}$ such that every point of $A$ lies on every hyperplane in $B$. We consider three cases: either $|A|\geq s^{2}\ell p^{t-2}$, $|B|\geq s^{2}\ell p^{t-2}$ or $\max\{|A|,|B|\}<s^{2}\ell p^{t-2}$. In all cases, we show that
$|A|\cdot|B|\leq s^{4}\ell p^{t-1}\leq 2s^{4}\varepsilon\ell p^{t}\leq 16s^{4}\varepsilon n.$
@@
To finish the proof, let $\mathcal{P}$ be a random subset of $\mathcal{P}_{0}$ containing $m$ points and $\mathcal{H}$ be a random subset of $\mathcal{H}_{0}$ containing $n$ hyperplanes. One should verify that $m\leq|\mathcal{P}_{0}|$ and $n\leq|\mathcal{H}_{0}|$, but this is clear since $|\mathcal{P}_{0}|,|\mathcal{H}_{0}|\geq\frac{1}{4}\ell p^{t}\geq n$ and $n\geq m$. We also have $\operatorname{rs}(\mathcal{P},\mathcal{H})\leq\operatorname{rs}(\mathcal{P}_{0},\mathcal{H}_{0})\leq(2s)^{4}\varepsilon n$. Furthermore, the expected number of incidences between $\mathcal{P}$ and $\mathcal{H}$ is
$\mathbb{E}\big{[}I(\mathcal{P},\mathcal{H})\big{]}=\frac{n}{n_{0}}\frac{m}{m_{0}}I(\mathcal{P}_{0},\mathcal{H}_{0})=\frac{1}{p}mn\geq\varepsilon mn.$
Thus, there exist subsets $\mathcal{P},\mathcal{H}$ of $\mathcal{P}_{0},\mathcal{H}_{0}$ with at least $\varepsilon mn$ incidences and $\operatorname{rs}(P,\mathcal{H})\leq(2s)^{4}\varepsilon n$. ∎
@@
Let $\varepsilon\in(0,1)$, and let $d$ and $m\leq n$ be positive integers which satisfy
$\varepsilon\leq\frac{1}{4}m^{-\frac{1}{d-1}}.$
Then there exists a prime power $q$, a set $\mathcal{P}$ of $m$ points and a set $\mathcal{H}$ of $n$ hyperplanes in $\mathbb{F}_{q}^{d}$ such that $I(\mathcal{P},\mathcal{H})\geq\varepsilon mn$ and $\operatorname{rs}(\mathcal{P},\mathcal{H})\leq(4s)^{4}\varepsilon n$, where $s$ is given by Lemma 3.14.
@@
More precisely, we apply Lemma 3.15 to find a set of $m$ points $\mathcal{P}_{0}$ and a set of $m$ hyperplanes $\mathcal{H}_{0}$ in $\mathbb{F}_{p}^{d-1}$ for some prime $p$ for which $I(\mathcal{P}_{0},\mathcal{H}_{0})\geq\varepsilon m^{2}$ and $\operatorname{rs}(\mathcal{P}_{0},\mathcal{H}_{0})\leq(2s)^{4}\varepsilon m$. The hyperplanes and points may be considered in the larger ambient space $\mathbb{F}_{q}^{d-1}$, for any prime power $q=p^{k}$. Passing to $\mathbb{F}_{q}^{d-1}$ does not change the incidence graph. In the rest of the proof, we fix $q$ to be a power of $p$ larger than $n/m$.
Now, for each $x\in\mathbb{F}_{q}^{d-1}$, let $x^{\prime}\in\mathbb{F}_{q}^{d}$ be the point whose first $d-1$ coordinates are the same as $x$, and the last coordinate is 00. Set $\mathcal{P}=\{x^{\prime}|x\in\mathcal{P}_{0}\}$. Furthermore, we fix a set $S=\{s_{1},\dots,s_{\ell}\}\subseteq\mathbb{F}_{q}$ of size $\ell=\lceil n/m\rceil$. Then, for every hyperplane $H\in\mathcal{H}_{0}$, if $H$ is given by the equation $\langle a,x\rangle=a_{1}x_{1}+\cdots+a_{d-1}x_{d-1}=b$, we define hyperplanes $H^{(1)},\dots,H^{(\ell)}$ in $\mathbb{F}_{q}^{d}$, where $H^{(i)}$ is given by the equations $a_{1}x_{1}+\cdots+a_{d-1}x_{d-1}+s_{i}x_{d}=b$. Then, we let $\mathcal{H}_{1}=\{H^{(i)}|i\in[\ell],H\in\mathcal{H}_{0}\}$.
The sizes of the newly constructed sets are $|\mathcal{P}|=m,|\mathcal{H}_{1}|=m\lceil\frac{n}{m}\rceil>n$. Furthermore, if $x\in\mathcal{P}_{0}$ is incident to $H\in\mathcal{H}_{0}$, then $x^{\prime}$ is incident to all of the hyperplanes $H^{(1)},\dots,H^{(\ell)}$. Thus, the number of incidences between $\mathcal{P}$ and $\mathcal{H}_{1}$ is $I(\mathcal{P},\mathcal{H}_{1})\geq\ell I(\mathcal{P}_{0},\mathcal{H}_{0})\geq\varepsilon\ell|\mathcal{P}_{0}||\mathcal{H}_{0}|=\varepsilon|\mathcal{P}||\mathcal{H}_{1}|$. Finally, the size of the largest complete bipartite graph is bounded by
$\operatorname{rs}(\mathcal{P},\mathcal{H}_{1})\leq\ell\cdot\operatorname{rs}(\mathcal{P}_{0},\mathcal{H}_{0})\leq\ell\cdot(2s)^{4}\varepsilon m\leq(4s)^{4}\varepsilon n.$
@@
Let $\varepsilon\in(0,1)$, and let $d$ and $m\geq n$ be positive integers which satisfy
$\varepsilon\leq\frac{1}{4}n^{-\frac{1}{d-1}}.$
Then there exists a prime power $q$, a set $\mathcal{P}$ of $m$ points and a set $\mathcal{H}$ of $n$ hyperplanes in $\mathbb{F}_{q}^{d}$ such that $I(\mathcal{P},\mathcal{H})\geq\varepsilon mn$ and $\operatorname{rs}(\mathcal{P},\mathcal{H})\leq(4s)^{4}\varepsilon m$, where $s$ is given by Lemma 3.14.
@@
The proof is verbatim the same as the proof of Lemma 3.16, the only difference being that we “blow up” points instead of hyperplanes. More precisely, we start from a set of $n$ points $\mathcal{P}_{0}$ and a set of $n$ hyperplanes $\mathcal{H}_{0}$ in $\mathbb{F}_{p}^{d-1}$ for which $I(\mathcal{P}_{0},\mathcal{H}_{0})\geq\varepsilon n^{2}$ and $\operatorname{rs}(\mathcal{P}_{0},\mathcal{H}_{0})\leq(2s)^{4}\varepsilon n$ and consider them in $\mathbb{F}_{q}^{d-1}$, where $q>m/n$.
Then, we fix a set $S=\{s_{1},\dots,s_{\ell}\}\subseteq\mathbb{F}_{q}$ of size $\ell=\lceil m/n\rceil$ and for each point $x\in\mathbb{F}_{q}^{d-1}$ we define $x^{(i)}\in\mathbb{F}_{q}^{d}$ be the point whose first $d-1$ coordinates are the same as $x$, and the last coordinate is $s_{i}$. Set $\mathcal{P}_{1}=\{x^{(i)}|x\in\mathcal{P}_{0},i\in[\ell]\}$. Then, for every hyperplane $H\in\mathcal{H}_{0}$, if $H$ is given by the equation $\langle a,x\rangle=a_{1}x_{1}+\cdots+a_{d-1}x_{d-1}=b$, we define the hyperplane $H^{\prime}$ in $\mathbb{F}_{q}^{d}$, given by the equation $a_{1}x_{1}+\cdots+a_{d-1}x_{d-1}+0\cdot x_{d}=b$. Then, we let =$\mathcal{H}=\{H^{\prime}|H\in\mathcal{H}_{0}\}$.
The sizes of the newly constructed sets are $|\mathcal{H}|=n,|\mathcal{P}_{1}|=n\lceil\frac{m}{n}\rceil>n$. Thus, by choosing a random subset $\mathcal{P}\subseteq\mathcal{P}_{1}$ and performing the same calculation as in the proof of Lemma 3.16 one arrives at the desired conclusion. ∎
@@
Let $\varepsilon\in(0,1)$, and let $d$ and $m,n$ be positive integers which satisfy
$4\varepsilon\geq n^{-\frac{1}{d-2}}\text{ and }4\varepsilon\geq m^{-\frac{1}{d-2}}.$
Then there exists a prime power $q$, a set $\mathcal{P}$ of $m$ points and a set $\mathcal{H}$ of $n$ hyperplanes in $\mathbb{F}_{q}^{d}$ such that $I(\mathcal{P},\mathcal{H})\geq\varepsilon mn$ and $\operatorname{rs}(\mathcal{P},\mathcal{H})\leq 2^{6}s^{4}(4\varepsilon)^{d-1}mn$, where $s$ is given by Lemma 3.14.
@@
The proof of this lemma is somewhat similar to the proofs of Lemmas 3.16 and 3.17, since we use the idea of “blowing-up” a lower-dimensional construction. However, in this case we use a $(d-2)$-dimensional configuration as a starting point and we “blow-up” both points and hyperplanes.
@@
As in the proof of Lemma 3.16, we consider $\mathcal{P}_{0},\mathcal{H}_{0}$ in $\mathbb{F}_{q}^{d-2}$ instead of $\mathbb{F}_{p}^{d-2}$, where $q$ is the power of $p$ satisfying $q>\max\{\frac{m}{N},\frac{n}{N}\}$. Furthermore, we fix two sets $S,T\subseteq\mathbb{F}_{q}$, such that $S=\{s_{1},\dots,s_{k}\}$ and $T=\{t_{1},\dots,t_{\ell}\}$, where $k=\lceil\frac{m}{N}\rceil,\ell=\lceil\frac{n}{N}\rceil$. To every point $x\in\mathcal{P}_{0}$, we associate $k$ points of $\mathbb{F}_{q}^{d}$ as follows. For $i\in[k]$, let $x^{(i)}\in\mathbb{F}_{q}^{d}$ be equal to $x$ on the first $d-2$ coordinates, equal to $s_{i}$ on the coordinate $d-1$, and equal to 00 on the last coordinate. Similarly, for a hyperplane $H\in\mathcal{H}_{0}$ given by the equation $a_{1}x_{1}+\cdots+a_{d-2}x_{d-2}=b$, we associate the hyperplanes $H^{(j)}$ in $\mathbb{F}_{q}^{d}$ for $j\in[\ell]$ defined by the equations $a_{1}x_{1}+\cdots+a_{d-2}x_{d-2}+0\cdot x_{d-1}+t_{j}x_{d}=b$. Finally, we set $\mathcal{P}_{1}=\{x^{(i)}|i\in[k],x\in\mathcal{P}_{0}\}$ and $\mathcal{H}_{1}=\{H^{(j)}|H\in\mathcal{H}_{0},j\in[\ell]\}$.
@@
Recalling that $N=(4\varepsilon)^{-(d-2)}$, we obtain $\operatorname{rs}(\mathcal{P}_{1},\mathcal{H}_{1})\leq 2^{6}s^{4}(4\varepsilon)^{d-1}mn$. Subsampling a random $m$ element subset $\mathcal{P}$ of $\mathcal{P}_{1}$, and a random $n$ element subset $\mathcal{H}$ of $\mathcal{H}_{1}$, completes the proof. ∎